OpenStudy (candyme):

What method would be best to solve this system of equations? (Addition, Subtraction, Subsitution) How would you solve it? Please include steps, because I want to learn how to do this.

OpenStudy (candyme):

\[2x+y=-5\]\[x+3y=25\]

OpenStudy (anonymous):

subtraction 2x+y=-5--------1 x+3y=25--------2 multiply the 2nd equation by 2 2(x+3y=25) --> 2x+6y=50 now subtract 2nd from 1st equation 2x+y=-5 -2x-6y=-50 ------------ -5y=-55 y=-55/-5=11 now plug in y value into one of the equation x+3(11)=25 x+33=25 x=25-33 x=-8 so x=-8 and y=11

OpenStudy (candyme):

Thank you :)